how many milligrams are present in 9.7748 x 10^22 molecules of chlorine gas at STP?
1
Expert's answer
2020-12-18T08:45:39-0500

Solution:

Molar mass of Cl2 70,914 g/mol

m = n×M

n = 9.7748×1022 / 6.0221×1024 = 0,016 mol Cl2

m = 0.016× 70.914= 1.135 g = 1 135 mg

Answer: mass of Cl2 is 1 135 mg.